# ChatWise: AI-Powered Engaging Conversations for Enhancing Senior Cognitive Wellbeing

### v1

@@ Abstract
Cognitive health in older adults presents a growing challenge. While conversational interventions show feasibility in improving cognitive wellness, human caregiver resources remain overburdened. AI-based methods have shown promise in providing conversational support, yet existing work is limited to implicit strategy while lacking multi-turn support tailored to seniors. We improve prior art with an LLM-driven chatbot named ChatWise  for older adults. It follows dual-level conversation reasoning at the inference phase to provide engaging companionship. ChatWise thrives in long-turn conversations, in contrast to conventional LLMs that primarily excel in short-turn exchanges. Grounded experiments show that ChatWise significantly enhances simulated users’ cognitive and emotional status, including those with Mild Cognitive Impairment.
ChatWise: AI-Powered Engaging Conversations
for Enhancing Senior Cognitive Wellbeing
Zhengbang Yang, Zhuangdi Zhu George Mason University, Fairfax, VA, USA {zyang30,zzhu24}@gmu.edu
@@ 1 Introduction
[…]
@@
In this work, we aim to draw on the advance of AI to provide older adults with engaging conversational support that improves their cognitive function and reduces social loneliness, serving as accessible alternatives to human companions. Previous clinical studies aimed at socially isolated older adults showed the existence of causal relationships between interviewer strategy and interviewee response  (Cao et al., 2021), revealing that conversation behavior can have a measurable influence on interlocutors, which inspired us to develop principled yet efficient methods that transform clinical insights into strategic dialogue design leveraging the advanced reasoning ability of LLMs.
@@
In response, we propose an LLM-driven chatbot named ChatWise. Unlike traditional chatbot driven with implicit conversational goals, ChatWise employs dual-layer conversation generation, which first derives categorized macro-level information, including user emotion states to suggest meta-conversational strategies, which then guides the micro-level utterance generation to improve both user engagement and cognitive outcomes over multi-turn dialogue interactions. Figure 1 overviews its development design. Our work provides multifold contributions:
  • ChatWise is a tuning-free yet principled framework for developing AI chatbots with enhanced multi-turn, daily-themed conversational support. Our dual-level conversation generation design integrates clinical insights into LLM reasoning and can be readily applied to various LLMs.
  • ChatWise follows grounded development and evaluation using digital twins (Hong et al., 2024), which are simulated users constructed by tuning LLMs on de-identified, real-world dialogues between seniors and professional caregivers of a clinical trial (I-CONECT, 2024). Interactions with digital twins demonstrated that ChatWise significantly enhances users’ engagement and cognitive status with multi-turn interactions, especially for users with Mild Cognitive Impairment (MCI) (Figure 2).
  • Our comparative studies demonstrated that integrating macro-level strategies into conversation generation is the key contributor to enhancing user engagement. Its gain on cognitive metrics remains consistent across different backbone LLMs. ChatWise’s advantages become more pronounced in longer conversations, in contrast to conventional LLMs that primarily excel in short-turn question-answer exchanges.
  • Our dialogue analysis revealed key empirical insights, including the pattern of predominant interaction strategies across users. While transient user emotional fluctuations were less tractable, multi-turn conversations with ChatWise have shown to notably improve user emotions over time. This highlights the value of long-term, principled conversational support in enhancing seniors’ cognitive and emotional well-being. We hope these findings can inspire future research, including RL-based LLM optimization.
Figure 1: Overview of ChatWise Development.
Figure 2: ChatWise notably improves multiple user cognitive metrics through 10-turn dialogues (Sec 4.2).
@@ 2 Related Work
AI-powered Chatbots for Seniors: Social interaction plays a crucial role in the cognitive health of seniors  (Patil and Braun, 2024; Fiori and Jager, 2012). AI-powered chatbots have emerged as a promising alternative  (Rodríguez-Martínez et al., 2023; Owan et al., 2023). Recent efforts span commercial products (sen, 2024; ell, 2024) and research prototypes focused on emotional support and audio assistance  (Yang et al., 2024; Liu et al., 2021b; sen, 2024; Ryu et al., 2020; Hong et al., 2024). Unlike prior work, our approach prioritizes user conversational engagement, which has been empirically shown to enhance both cognitive and emotional status.
Conversational Strategies: Liu et al. (2021a) curated a dataset with annotated strategies, demonstrating the effectiveness of Helping Skills Theory Hill (2020) in providing emotional support. Yuan et al. (2023) examined the causal relationships between dialogue acts (DAs) and participants’ emotional states in a clinical trial (I-CONECT, 2024), emphasizing the impact of strategic interventions in tele-mediated dialogues. Seo et al. (2021) identified key strategies for improving child patient-provider communication through semi-structured interviews. Few works have systematically integrated these strategies into the reasoning flow of AI chatbots. Our approach fills this gap by embedding structured conversational strategies for automatically enhancing user engagement.
Multi-turn chatbot exploration: Recent advances in AI dialogue systems predominantly focused on short-turn interactions  (Owan et al., 2023; Dam et al., 2024). While some pioneering research has explored multi-turn optimization through Reinforcement Learning (RL) approaches for LLMs (Verma et al., 2022; Zhou et al., 2024; Abdulhai et al., 2023; Gao et al., 2025), these methods were not specifically designed for supporting senior dialogue engagement. Our approach is tuning-free yet compatible with RL-driven methods, enabling future extensions through hierarchical RL to further enhance multi-turn conversation engagement.
@@ 3 ChatWise Design
ChatWise employs a dual-level framework with a strategy provider $\pi_{s}$ that induces macro-level actions, and an utterance generator $\pi_{u}$ for deriving utterances based on suggested strategies (Figure 1).
Strategy Candidates: To construct a strategy candidate set, we extract Dialogue Acts (DAs) from real telehealth clinical trials Yuan et al. (2023), where DAs serve as atomic communicative units that convey distinct conversational intentions Searle (1976). We further integrate these with strategies from prior emotional support dataset (Liu et al., 2021a) to form a comprehensive set. We refine each strategy with definitions and examples as in-context prompts to enhance $\pi_{s}$ reasoning.
From Strategy to Utterance: $\pi_{s}$ reasons over the dialogue history $\tau$ to capture user’s emotion $e$ and derive one to two strategies, $\{s_{i}\}_{i\leq 2}\sim\pi_{s}(\tau,e)$. Our rationale follows counseling study (Zhang and Danescu-Niculescu-Mizil, 2020) that dialogue intentions can be forward, e.g. initiating new topics via an open question, or backward, e.g. responding with acknowledgment. $\pi_{s}$ is prompted to flexibly employ one or both as needed. Selected strategies, user’s current emotion, and conversation history serve as inputs for $\pi_{u}$ to generate utterances. To ensure natural flow, $\pi_{u}$ will first improvise a few rounds before following suggested strategies. Drawing on clinical guidance, $\pi_{u}$ encourages users to choose topics rather than imposing them.
4 Empirical Evaluation
Below we summarize experiment configurations, metrics, and main findings of ChatWise. More experimental details are available in Appendix A.1 and A.2.
4.1 Conversation Generation

### v2

# ChatWise: A Strategy-Guided Chatbot for Enhancing Cognitive Support in Older Adults

@@ Abstract
Cognitive health in older adults presents a growing challenge. Although conversational interventions show feasibility in improving cognitive wellness, human caregiver resources remain overloaded. AI-based chatbots have shown promise, yet existing work is often limited to implicit strategies or heavily depends on training and label resources. In response, we propose a strategy-guided AI chatbot named ChatWise  that follows a dual-level conversation reasoning framework. It integrates macro-level strategy planning and micro-level utterance generation to enable engaging, multi-turn dialogue tailored to older adults. Empirical results show that ChatWise closely aligns with professional human caregiver behaviors in offline evaluation using real clinic data, and achieves positive user cognitive and emotional responses in interactive simulations with digital twins, which significantly outperforms AI baselines that follow implicit conversation generation.
ChatWise: A Strategy-Guided Chatbot for Enhancing Cognitive Support in Older Adults
Zhengbang Yang1, Junyuan Hong2, Yijiang Pang3, Jiayu Zhou4, Zhuangdi Zhu1 1 George Mason University, Fairfax, VA, USA 2 University of Texas at Austin, Austin, TX, USA 3 Michigan State University, East Lansing, MI, USA 4 University of Michigan, Ann Arbor, MI, USA {zyang30,zzhu24}@gmu.edu
@@ 1 Introduction
The cognitive well-being of the elderly population is a pressing social concern, as evidenced by the prevalence of cognitive disorders within this population, often exacerbated by loneliness and isolation Nicholson (2012); Teo et al. (2023). According to WHO, approximately $14\%$ of adults aged 60 and over experience mental health disorders, projected to affect 2.1 billion individuals by 2050 Organization (2024). Compared with other age groups, older adults are more vulnerable due to age-related changes in cognitive reserves  (Salthouse, 2009) and reduced social connections  (Nicholson, 2012; Teo et al., 2023). Such impact extends beyond older individuals to families and society, resulting in a reduced life quality and increased medical burden. In the meantime, modest delays in cognitive decline can significantly reduce dementia prevalence, and addressing social isolation could prevent 4% of dementia cases Livingston et al. (2020). Interventions through guided conversations have shown efficacy in reducing loneliness and mitigating cognitive decline  (Yu et al., 2021, 2022, 2023; Fiori and Jager, 2012). However, effective intervention requires sustained interaction and monitoring, which is limited by the availability of human companions, leading to inconsistent access or effectiveness.
Advances in artificial intelligence (AI) particularly Large Language Models (LLMs) have shown promise in augmenting human expertise with conversational support. Existing efforts range from sentiment-based chatbots Liu et al. (2021); sen (2024); Ryu et al. (2020); Tsai et al. (2021); You et al. (2023); Meyer et al. (2020) to audio assistants Yang et al. (2024); Hong et al. (2024), ascribed to mature text-audio transformation techniques Van Den Oord et al. (2016); Ren et al. (2019). However, these systems primarily default to interactions with implicit goals, which may fail to drive engaging conversations that necessitate strategic multi-turn interactions tailored to older adults. Some work has explored fine-tuning LLMs on domain-specific datasets to adapt to elderly care Sun et al. (2025), which requires extensive labeled data and resources, while they may struggle to generalize across conversational contexts. In contrast, we take an orthogonal approach by focusing on the inherent reasoning capabilities of LLMs through in-context learning, which offers more resource efficiency and adaptability.
@@
Our goal is to provide AI-powered, engaging conversational support for older adults that serves as an accessible complement to human companions, with the aspiration to improve their cognitive function and reduce social loneliness. Previous clinical studies aimed at socially isolated older adults showed the existence of causal relationships between interviewer strategy and interviewee response  (Cao et al., 2021), which revealed that conversation behavior can have a measurable influence on interlocutors, and thus inspired us to develop principled yet efficient methods that transform clinical insights into dialogue design by leveraging the advanced reasoning ability of LLMs.
@@
In response, we propose an LLM-driven chatbot named ChatWise. It employs dual-layer conversation generation that first derives categorized macro-level information to suggest meta-conversational strategies, which then guides the micro-level utterance generation to improve both user engagement and cognitive outcomes over multi-turn dialogue interactions. ChatWise is evaluated using both real, de-identified real-world dialogues between older adults and professional human caregivers of a clinical trial (I-CONECT, 2024), and synthetic interactive conversations using simulated users (i.e. digital twins)  (Hong et al., 2024) modeled on such data.
Our work provides multifold contributions: (i) We introduce a clinically grounded chatbot that highly aligns with the behavior of professional human caregivers to older adults (Sec 4.1). Figure 1 overviews its alignment performance. It also empirically enhances simulated users’ engagement and cognitive status, which significantly outperforms baseline chatbots (Sec 4.2). (ii) ChatWise follows a tuning-free, in-context learning framework for daily conversational support. Comparative studies demonstrated that providing macro-level strategies to guide conversation generation is the key contributor to enhancing user engagement. (iii) Its dual-level policy design can be readily applied to various LLMs, with consistent user cognitive gains across different backbone LLMs.
Figure 1: Alignment between conversational strategies of AI chatbots and human professionals across diverse participant data. Our design shows closer alignment with caregiver behavior in real conversation contexts compared to baseline chatbots. See Sec 4 for full results.
@@ 2 Related Work
AI-powered Chatbots for Older Adults
have shown feasibility in preventing or detecting the cognitive issues to assist human professionals. Recent efforts span commercial products (sen, 2024; ell, 2024) and research prototypes focused on emotional support and audio assistance. (Sun et al., 2025) performed supervised fine-tuning on LLMs to enhance their performance on specialized nursing and elderly care tasks.  Yang et al. (2024) introduced an LLM-based voice assistant designed to bridge communication between older adults and their healthcare providers. Liu et al. (2021) built an annotated dataset to tune LLMs for emotional support tasks. Ryu et al. (2020) developed a chatbot for the mental health of the elderly. Hong et al. (2024) constructed digital twins of the elderly. Unlike prior work, our approach prioritizes user conversational engagement through in-context reasoning, which has been empirically shown to enhance user cognitive status.
Dialogue Systems for Mental Health or Cognitive Stimulation:
Recent studies have leveraged LLMs for the augmentation of emotional support conversations through generated dialogue data  (Zheng et al., 2023; Jiang et al., 2023).  Xygkou et al. (2024) assessed the acceptability of GPT-4-based conversational agents among people with dementia (PwD), and highlighted the importance of design considerations for this sensitive population. Similarly,  Favela et al. (2023) implemented a conversational robot embedded with ChatGPTs for reminiscence therapy, with authors as simulated users, which showcased the potential of generative AI to support cognitive stimulation through interactive storytelling. Distinct from these works, our research introduces a dual-level dialogue approach combined with comprehensive evaluation methodologies to provide a finer-grained understanding of conversational engagement.
Conversational Strategies:
Liu et al. (2021) curated a dataset with annotated strategies, demonstrating the effectiveness of Helping Skills Theory Hill (2020) in providing emotional support. Yuan et al. (2023) examined the causal relationships between dialogue acts (DAs) and participants’ emotional states in a clinical trial (I-CONECT, 2024), emphasizing the impact of strategic interventions in tele-mediated dialogues. Seo et al. (2021) identified key strategies for improving child patient-provider communication through semi-structured interviews. However, few works have systematically integrated these strategies into the reasoning flow of AI chatbots. Our approach fills this gap by contextualizing structured conversational strategies for automatically enhancing user engagement.
Multi-Turn Chatbot Exploration:
Recent advances in AI dialogue systems predominantly focused on short-turn interactions  (Owan et al., 2023; Dam et al., 2024; Gao et al., 2025), with relatively limited attention to the challenges of multi-turn exchanges. While some pioneering research has explored multi-turn optimization through Reinforcement Learning (RL) approaches for LLMs (Verma et al., 2022; Zhou et al., 2024; Abdulhai et al., 2023; Gao et al., 2025), these methods were not specifically designed for supporting senior dialogue engagement. Our inference-based method introduces a strategy-compatible framework that aligns with the principles of RL-driven optimization, which can enable future extension to further enhance multi-turn conversation engagement.
Figure 2: ChatWise  employs a dual-level policy design that contextualizes a clinically-derived strategy pool to generate macro-level actions, which then guide micro-level utterance generation. It is iteratively developed using digital twins as simulated users. Prior to a future real user study, real and simulated data analysis show that ChatWise closely aligns with professional human caregivers’ behavior, leading to cognitive gains of simulated users accumulated over multi-turn interactions.
@@ 3 ChatWise Design
ChatWise encompasses a hierarchical framework that integrates the insights of traditional clinical studies to drive LLM reasoning. It features a dual-level design with a strategy provider $\pi_{s}$ that generates macro-level conversation strategies, which guides a utterance generator $\pi_{u}$ to produce micro-level responses. In this paper, we interchangeably refer to ChatWise as the interviewer or the moderator, a role commonly defined in communication studies that guides and facilitates the conversation  (Taboada, 2006),
For clarity, we denote a multi-turn conversation as $\tau=\{\mu_{0},u_{0},x_{1},u_{1},x_{2},u_{2},\cdots\}$ which comprises interviewee (user) utterances $x_{t}$ and interviewer (ChatWise) utterances $u_{t}$ at time step $t$, with a conversation starting point $\mu_{0}$. $s_{t}=\{\mu_{0},u_{0},\cdots,x_{t}\}$ denotes the historical conversation up to timestep $t$. We then formulate ChatWise as a dual policy $\pi=\{\pi_{s},\pi_{u}\}$, which maps the conversational state $s_{t}$ to the next utterance $u_{t}$ provided to the user, conditioned on meta information $z$ given to the strategy provider $\pi_{s}$:
$u_{t}\sim\pi(\cdot|s_{t},z)\equiv\pi_{u}(\cdot|\pi_{s}(s_{t}|z)).$
Below, we illustrate the process of this hierarchical conversation generation.
3.1 Strategy Pool
To first develop clinically grounded strategies, we establish a strategy pool containing dialogue acts (DA) curated from clinical intervention studies involving older adults. These DAs, such as asking an open question (e.g. “What is your favorite movie”) or showing empathy (e.g.“That sounds really meaningful”), serve as macro-level actions and atomic communicative units that convey distinct conversational intentions Searle (1976), which have also shown causal relationships with participant emotions in conversations Cao et al. (2021). Specifically, we extract DAs that exist in conversations from a real tele-health clinical trial Yuan et al. (2023), which represents a subset of the complete DAMSL taxonomy Allen and Core (1997); Searle (1976). We further augment this DA with strategies from prior emotional support dialogue dataset (Liu et al., 2021) to form a comprehensive strategy pool $\mathcal{A}$. We then treat each strategy serves as a macro-level action candidate $a\in\mathcal{A}$ and refine it with a definition and an example to serve as in-context information $z$ for enhancing the reasoning of strategy $\pi_{s}$ (See Appendix A.5 for a complete DA set).
3.2 Emotion Annotation
To enhance context-awareness during conversations and support post-conversation analysis, the strategy provider $\pi_{s}$ is additionally prompted to infer the user’s current emotion based on historical conversations $s_{t}$. At timestep $t$ of the conversation, $\pi_{s}$ will categorize user emotion into one of the five classes: joy, neutral, sadness, anger, and surprise, given a calibrated system prompt. Our comparative study shows that ChatWise can lead to detectable positive emotion changes over multi-turn conversations, which aligns with conversations provided by human professionals (Sec 4.5).
3.3 From Strategy to Utterance Generation
The strategy provider $\pi_{s}$ processes the dialogue history $s_{t}$, contextualized by the strategy pool information $z$ to decide the most appropriate strategies for continuing a conversation. We limit each interviewer’s utterance to contain one or two strategies, i.e. $\mathbf{a}_{t}=\{a_{t}^{i}\}_{i\leq 2}\sim\pi_{s}(\cdot|s_{t},z)$. Our rationale follows counseling studies (Zhang and Danescu-Niculescu-Mizil, 2020) that dialogue intentions can be forward, e.g.  initiating new topics via an open question, or backward, e.g.  responding with acknowledgment, or both to transit between intentions or topics. Thus, $\pi_{s}$ is prompted to comply with one of the three conditions: (i) a forward strategy, (ii) a backward strategy, or (iii) a backward strategy followed by a forward strategy.
The selected strategies $\mathbf{a}_{t}$, and optionally, the user’s current emotion label $e_{t}$, and the conversation history $s_{t}$ are used as inputs to the utterance generator $\pi_{u}$. Our ablation study showed that providing meta-level strategies has a dominant effect over providing emotion information alone for conversational engagement (Sec 4.3). To ensure natural flow, $\pi_{u}$ will first improvise a few rounds as warm-ups before adopting the suggested strategies. We also draw on clinical guidance to let $\pi_{u}$ encourage users to choose topics rather than imposing them. Figure 2 overviews this design.
4 Experiments
We conducted various experiments focusing on answering the following questions:
Q1. Does ChatWise’s behavior align with human professionals when responding to real conversation context sampled from clinic trials?
Q2. How does ChatWise compare to baseline AI alternatives in terms of multifaceted conversational engagement metrics?
Q3. How do factors such as the conversation turns and different user characteristics influence ChatWise’s performance?
To address these questions, we evaluated ChatWise on two complementary scenarios: we first conducted offline evaluations using data from a real clinical trial (Sec 4.1), then we performed interactive experiments, in which conversations are stochastically generated between ChatWise and digital twins as simulated users (Sec 4.2). Below, we summarize the configurations, metrics, and main results of each experimental setting and defer more details to the Appendix.
4.1 Offline Evaluation on Real Clinical Data
4.1.1 Data Preprocessing
We extracted dialogues from the I-CONECT clinic trials I-CONECT (2024), with each in text format converted from video conversations. Short dialogues below 40 turns were excluded, where a turn represents an uninterrupted, continuous utterance by a single speaker. Two dialogue sets were subsampled and de-identified: (1) 150 randomly sampled dialogues to assess the overall strategy alignment, and (2) one dialogue per week from 7 randomly selected participants to assess ChatWise’s alignment robustness to different participants and time periods over the clinical trial.
4.1.2 Offline Evaluation Metric
We define Strategy Match Percentage (SMP) to quantify the alignment between the conversational strategies generated by a chatbot and those generated by a human professional caregiver from the I-CONECT clinic data. Let $\mathbf{a}_{t}$ denote the set of caregiver-provided strategies annotated from a real conversation $\tau$ at turn (timestep) $t$, and $\mathbf{a}^{\prime}_{t}$ the strategies from a simulated moderator (ChatWise or utterance generator), conditioned on the same real context $s_{t}$. Let $\mathbb{I}(a^{\prime}_{t}\in\mathbf{a}_{t})$ be an indicator function that returns 1 if a chatbot-generated strategy $a^{\prime}_{t}$ matches one of the human strategies $\mathbf{a_{t}}$, and 0 otherwise. The SMP is computed as following:
$\text{SMP}_{t}(\mathbf{a}_{t},\mathbf{a}^{\prime}_{t})=\frac{1}{|\mathbf{a}_{t}|}\sum_{a^{\prime}_{t}\in\mathbf{a}^{\prime}_{t}}\mathbb{I}(a^{\prime}_{t}\in\mathbf{a}_{t}).$
Figure 3: Comparison of moderator strategies: Given real dialogue contexts, responses from a baseline (generated), ChatWise (generated), and the human moderator (real) are compared w.r.t. underlying strategies. A real sample is shown here while omitting participant data.
4.1.3 Strategy Label Annotation
Given a real dialogue denoting as $\tau=\{\mu_{0},u_{0},x_{1},u_{1},x_{2},u_{2},\cdots\}$ with caregiver (moderator) utterance $u_{t}$ and participant utterance $x_{t}$ at turn $t$, we leveraged the macro-level strategy definition in Sec 3.1 to prompt an LLM as annotator (GPT-4o) and detect strategies $\mathbf{a}_{t}$ behind $u_{t}$ as the golden label. Denoting $s_{t}=\{\mu_{0},u_{0},x_{1},\cdots,x_{t}\}$ as the real conversation context, we feed $s_{t}$ to a baseline chatbot to get utterance action $\tilde{u}_{t}$. The baseline is identical to the utterance generator $\pi_{u}$ of ChatWise yet does not receive strategy guidance from $\pi_{s}$. Similarly, we get strategy labels $\tilde{\mathbf{a}}_{t}$ behind $\tilde{u}_{t}$ using the above annotator. We then input $s_{t}$ to ChatWise’s strategy provider $\pi_{s}$ to get proposed strategy $\mathbf{a}^{\prime}_{t}\sim\pi_{s}(s_{t})$. We report the strategy alignment of $\text{SMP}_{t}(\mathbf{a}_{t},\mathbf{a}^{\prime}_{t})$ (w/ ChatWise) and $\text{SMP}_{t}(\mathbf{a}_{t},\tilde{\mathbf{a}}_{t})$ (w/o ChatWise) over different turns $t$ and average the result across dialogues. Figure 3 illustrates the process of collecting each moderator’s response based on real conversation history.
4.1.4 Overall Strategy Alignment
As shown in Figure 5, which averages results over 150 dialogues, ChatWise consistently achieves an SMP close to 1.0, indicating a strong alignment with the strategies employed by real caregivers throughout the conversation. In contrast, the baseline chatbot deviates from human behavior when lacking action guidance. Both chatbots used GPT-4o as the utterance generator $\pi_{u}$, and o3-mini was used for $\pi_{s}$ in ChatWise. We do not employ micro-level utterance similarity, as the I-CONECT interventions involve spontaneous daily conversations, rather than fixed question answering, unlike domains in programming or math. Whereas, on a macro level, ChatWise mostly mirrors the strategic choices of human professionals who are well trained to follow clinical protocols.
To illustrate how our method behaves differently from the baseline, we collect the real dialogue context and different moderator responses where $\mathbf{a_{t}}=\mathbf{a}^{\prime}_{t}\neq\tilde{\mathbf{a}_{t}}.$ and show the discrepancy heatmap. Raw dialogue content is omitted to comply with the data usage agreement. Results in Figure 4 demonstrate that while “open question” is a commonly dominant strategy (forward DA) for both human and chatbot moderators, the primary strategy difference lies in how they lead into the question (e.g. “acknowledge”, “restatement”), which may in turn influence the specific question ultimately asked.
Figure 4: Strategy distribution for dialogue samples where $\mathbf{a_{t}}=\mathbf{a}^{\prime}_{t}\neq\tilde{\mathbf{a}_{t}}.$ The vertical axis indicates ground-truth strategies $\mathbf{a_{t}}$ extracted from human professional utterances, and the horizontal axis denotes strategies $\tilde{\mathbf{a}}_{t}$ detected from the baseline chatbot.
Figure 5: SMP between human caregiver and chatbots on real conversation data, where w/o ChatWise denotes the baseline without strategy guidance. Our method strongly aligns with human caregiver behavior.
Robustness to Participant Heterogeneity:
Figure 6 presents the SMP of ChatWise across dialogue turns among participants. ChatWise consistently outperforms the baseline and maintains notably higher SMP. These results demonstrate the robust alignment of ChatWise  with human caregiver strategies across diverse participants and their interaction contexts, which implies its potential for personalized dialogue systems.
Figure 6: SMP broken down by participants over 40 conversation turns, where each color indicates a specific person. Solid lines represent ChatWise, and dashed lines denote the baseline. ChatWise consistently aligns to human caregivers given different participant contexts.
Alignment Consistency Over Timeline:
Figure 7 presents the SMP over time, where dialogues were sampled from different weeks over the clinical intervention. ChatWise maintains consistently high SMP (near 1.0) across all timesteps, where the baseline shows lower and unstable performance. This indicates ChatWise’s temporal robustness and its potential to facilitate the development of long-term cognitively supportive dialogues.
Figure 7: Strategy alignment of ChatWise is consistent as real conversations progress over weeks.
4.2 Evaluation with Digital Twins
4.2.1 Conversation Generation